OpenStudy (loading):

A system of equations is shown below: 5x - 5y = 10 3x - 2y = 2 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (loading):

@mathmath333

OpenStudy (loading):

@TrojanPoem

OpenStudy (loading):

Can you help?

OpenStudy (loading):

@Opcode

OpenStudy (trojanpoem):

The questions asks you to take the any of the given equations for ex : 3x - 2y = 2 get a multiple of it and sum it to 5x - 5y = 10

OpenStudy (trojanpoem):

so take 3x - 2y = 2 ( * 2) 6x - 4y = 4 + 5x - 5y = 10 11x - 9 y = 14 new system of equations is 11x - 9y = 14 5x - 5y = 10

OpenStudy (loading):

A multiple of what number?

OpenStudy (loading):

I fairly confused.

OpenStudy (trojanpoem):

Any number 2 , 3, 4, 100000

OpenStudy (trojanpoem):

zero -> will destroy your equation 1 -> won't matter

OpenStudy (loading):

So, (3x - 2y = 2)*2 6x - 4y = 4 Is that what I should write.

OpenStudy (trojanpoem):

No it says Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other you got a multiple of the other (6x - 4y = 4) now it asks for the sum of that multiple (6x - 4y = 4) and the other equation ( 5x - 5y = 10)

OpenStudy (loading):

So, (3x - 2y = 2)*2 6x - 4y = 4 6x - 4y = 4 + 5x - 5y = 10 11x - 9y = 14 5x - 5y = 10 ^ Is that what I should write?

OpenStudy (loading):

@TrojanPoem

OpenStudy (loading):

@Kainui

OpenStudy (loading):

@Somy

OpenStudy (loading):

@tHe_FiZiCx99

OpenStudy (loading):

@Opcode

OpenStudy (loading):

@Lovelarap

OpenStudy (loading):

@hartnn

OpenStudy (trojanpoem):

yeah

OpenStudy (loading):

(3x - 2y = 2)*2 6x - 4y = 4 6x - 4y = 4 + 5x - 5y = 10 11x - 9y = 14 5x - 5y = 10 this is correct for A

OpenStudy (loading):

Now I need to solve for x and y? Correct"?

OpenStudy (anonymous):

if i'm not wrong yes

OpenStudy (loading):

Ok thanks everyone!